# Conceptualization, methodology, software, formal analysis, investigation, visualization, validation, writing – original draft

### v2

@@ 1 Introduction
Laser-scanning data can vary based on different characteristics, such as the platform on which the sensors are installed or the format in which the data are recorded. In terms of the platform, along with Terrestrial Laser Scanning or Mobile Laser Scanning, we have Airborne Laser Scanning (ALSD), Unmanned Laser Scanning (ULSD), and Space Laser Scanning (SLSFW) platforms. Owing to their cost-efficiency, coverage, and scalability, ALS and ULS are the most suitable options for mapping biomass at the local or regional scale. In contrast, SLS allows the mapping of biomass both nationally and globally (Luo et al., 2019). Likewise, laser-scanning data can be separated into discrete- and full-waveform data according to the format of the recorded information. Discrete data record the elements whose intensity value is the highest in the shape of a cloud of 3D points (such as treetops, branches, trunks, and ground). In contrast, full-waveform data record the entire waveform of the laser pulse that passes through the different vertical layers of the forest (citetcrespo2020processing. Discrete laser scanning has proven to be appropriate for most current applications. Nevertheless, full-waveform laser scanning can characterize the intermediate and lower layers of vegetation in a precise and detailed manner (Crespo-Peremarch, 2020). Despite this precision and detailed characteristics, there is still debate surrounding the type of laser scanning and the most appropriate data format for estimating AGB in tropical forests (Sanchez-Azofeifa et al., 2017).
To achieve model parsimony in explaining the AGB, we strive to employ the fewest possible variables, tackling the challenge of laser-scanning discrete point clouds potentially generating an excess of variables. This parsimonious approach was previously applied to GEDI data by Duncanson et al. (2022). In addition, Generalized Additive Models (GAMs) are employed to classify model variables by their significance in biomass prediction, a method previously utilized in a model by Lee et al. (2023) and Nandlall and Millard (2020). Once the candidate variables for the model are selected, the hyperparameters for the SVM regression are explored (Belete and Manjaiah, 2021). SVM regression can be employed with various kernel options, including linear, polynomial, radial basis function (RBF), and sigmoid (Yang and Shami, 2020). The hyperparameters can be determined, and the final set of variables for the regression model can be defined using the leave-one-out cross-validation method (Kudo et al., 2023). After fitting the regression model, it is crucial to analyze the errors to prevent underfitting or overfitting citepRamasubramanian. The evolution of errors in each iteration of the regression can be assessed to verify the absence of overfitting and gauge the overall quality of the final model.
@@ 2.1.1 Forest inventory of permanent plots
AGB was estimated for each tropical tree species using Equation 1 from Calvo-Rodriguez et al. (2021):
@@ 2.5 Regression models
The modeling workflow (fig.˜4) encompassed the following:
@@ 3.1 Exploratory Analysis, Variable Selection, and Predictive Significance
The research used 2021 forest inventory data on tree height and diameter at breast height (DBH) to model the above-ground biomass ($AGB$) using equation (1)). A standard Discrete Approach (DA) was applied: the total $AGB$ at each plot was calculated by summing the biomass of all individual trees, and the mean tree-level $AGB$ was then derived by dividing the total plot-level $AGB$ by the number of trees. Figure˜8 shows the values obtained from the inventory of DBH, height, and $AGB$ for the plots in this study.
@@ 3.2 Correlation analysis among variables
The correlation levels and significance of the values from the exploratory analysis are shown in Figures figs.˜9, 10 and 11. These correlograms display the accepted forest metrics for $ALS_{D}$, $ULS_{D}$, and $SLS_{FW}$. Significant values are indicated by white “X” marks, and the correlations among the accepted variables are shown in the diagrams.
@@
Figure˜12(a) illustrates the variable importance for $AGB_{t}$ using the $ALS_{D}$ dataset. After applying a threshold of 0.5, five variables were retained: Elev.minimum, Elev.L3, Elev.MAD.mode, Elev.mode, and Elev.MAD.median, while one variable fell below this threshold and was removed. These five predictors demonstrated the strongest influence on the $AGB_{t}$ regression. For $AGB_{m}$, the relative importance of the variables is shown in Figure˜12(b). Three variables, Elev.mode, Elev.MAD.mode, and Elev.L3 surpassed the 0.5 threshold and were chosen for the $AGB_{m}$ model.
@@
Figure˜12(c) displays the variable importance for predicting $AGB_{t}$ with the $ULS_{D}$ dataset. Initially, applying the 0.5 cutoff left only one variable (Elev.skewness). To avoid underfitting, the next most influential variable (Elev.minimum) was also included, resulting in the selection of two predictors: Elev.skewness and Elev.minimum. For $AGB_{m}$ (Figure˜12(d)), three variables met the 0.5 importance threshold: Elev.maximum, Elev.minimum, and Elev.MAD.mode. Therefore, these were chosen as the final set of predictors for the $ULS_{D}$-based $AGB_{m}$ model.
@@
Figure˜12(e) shows the importance of the predictors for the simulated $SLS_{FW}$ dataset in estimating $AGB_{t}$. Strictly applying the 0.5 cutoff led to no variables meeting the criterion. Consequently, the selection rule was adjusted to include the three most influential predictors: PointDense, gLAI0t10, and lat. Finally, Figure˜12(f) depicts the importance of the variables for $AGB_{m}$ estimation using the $SLS_{FW}$ data. Here, five variables, lat, pointDense, inflGround, blairSense, and maxGround, emerged as the most significant and were thus retained for the final model.
@@ 3.4 Regression analysis
The predictive capacities of the selected models were evaluated and compared with field measurements. Figure˜13 shows that the models achieved high accuracy in predicting $AGB$.
@@ 3.4.2 Results from the SVM approach
A grid search method was employed to identify the optimal hyperparameters that minimized the estimation error. Table˜4 illustrates the error progression during this optimization phase, depicting how the adjustments to the hyperparameters impacted the model’s precision.
@@
The analysis in Table 6 highlights the performance of SVM-based regression models across the three LiDAR platforms: ALS, ULS, and SLS. For ALS, two combinations of input variables yielded strong results. For AGBt, the best-performing variables were Elev.minimum, Elev.L3, and Elev.MAD.mode, whereas for AGBm, the most effective variables were Elev.mode and Elev.MAD.median. The estimation error (SVMr error) for AGBt was 18.0%, while for AGBm it was 13.1%, indicating that mean biomass is easier to estimate using ALS data. The optimal hyperparameters for ALS were $c=4$, $\sigma=0.25$ for AGBt, and $c=5$, $\sigma=2$ for AGBm.
@@
SVM regression was then conducted using the identified hyperparameters, and the resulting estimates were compared against the forest inventory for both AGBt and AGBm. Figure˜15 provides a synthesized view of the percentage errors, highlighting both the discrepancies and the accuracy of the model predictions.
@@ 3.5 Comparison between OLS and SVM AGB Estimates
The Mean Absolute Error (MAE) heatmap provides valuable insights into the variability of biomass estimates and their associated errors. As shown in Table 7, the differences in biomass estimates are directly correlated with the magnitude of error observed in the measurements or calculations. The upper diagonal of the table quantifies these errors, offering a clear view of how inaccuracies can propagate and lead to discrepancies in the final biomass estimates. The lower section of table˜7 reports the statistical significance of these variations, indicating the confidence levels associated with each comparison.
@@
As shown in Table˜8, the maximum percentage error observed in AGB modeling reached 92.7% for AGBt on plot L2.2 when utilizing ULSD data. Similarly, the most substantial modeling error for AGBm was 46.4%, recorded on the L2.2 plot using SLSFW data. In a broader context, the estimation errors for AGB were higher in L2.1 and L2.2 late-stage plots compared to L1.2, L1.1, and early and intermediate plots (E and I). The differences found in the AGB regression when using ALSD or ULSD data were relatively low, and the differences between the mean errors of ALSD and ULSD with the SLSFW data only reached 0.92% of the mean errors for AGBm and 0.68% for AGBm.
@@
The accuracy of AGB modeling using SVM exceeded 81% (ALSD: 82.0%, ULSD: 81.4%, and SLSFW: 82.9%) (table˜8), whereas the accuracy of OLS modeling was approximately 60%.
@@ 3.6 Results of AGB estimates for Forest Inventory, ALSD, ULSD, and SLSFW
Figure˜17 demonstrates the relationship between predicted above-ground biomass (AGB) and inventory-based AGB, expressed in Mg ha-1, across three LiDAR systems (ALSD, ULSD, and SLSFW) and different forest successional stages (Early, Intermediate, and Late). The x-axis represents the inventory-derived AGB values, considered the ground truth, whereas the y-axis displays the AGB predicted using the three LiDAR systems. A dashed diagonal line (1:1 line) is included to represent perfect agreement between the predicted and inventory-based AGB. The closer the points and trend lines are to this diagonal, the greater the prediction accuracy.

### v4

@@ 1 Introduction
Laser-scanning data can vary based on different characteristics, such as the platform on which the sensors are installed or the format in which the data are recorded. In terms of the platform, along with Terrestrial Laser Scanning or Mobile Laser Scanning, we have Airborne Laser Scanning (ALSD), Unmanned Laser Scanning (ULSD), and Space Laser Scanning (SLSFW) platforms. Owing to their cost-efficiency, coverage, and scalability, ALS and ULS are the most suitable options for mapping biomass at the local or regional scale. In contrast, SLS allows the mapping of biomass both nationally and globally (Luo et al., 2019). Likewise, laser-scanning data can be separated into discrete- and full-waveform data according to the format of the recorded information. Discrete data record the elements whose intensity value is the highest in the shape of a cloud of 3D points (such as treetops, branches, trunks, and ground). In contrast, full-waveform data record the entire waveform of the laser pulse that passes through the different vertical layers of the forest (Crespo-Peremarch, 2020). Discrete laser scanning has proven to be appropriate for most current applications. Nevertheless, full-waveform laser scanning can characterize the intermediate and lower layers of vegetation in a precise and detailed manner (Crespo-Peremarch, 2020). Despite this precision and detailed characteristics, there is still debate surrounding the type of laser scanning and the most appropriate data format for estimating AGB in tropical forests (Sanchez-Azofeifa et al., 2017).
To achieve model parsimony in explaining the AGB, we strive to employ the fewest possible variables, tackling the challenge of laser-scanning discrete point clouds potentially generating an excess of variables. This parsimonious approach was previously applied to GEDI data by Duncanson et al. (2022). In addition, Generalized Additive Models (GAMs) are employed to classify model variables by their significance in biomass prediction, a method previously utilized in a model by Lee et al. (2023) and Nandlall and Millard (2020). Once the candidate variables for the model are selected, the hyperparameters for the SVM regression are explored (Belete and Manjaiah, 2021). SVM regression can be employed with various kernel options, including linear, polynomial, radial basis function (RBF), and sigmoid (Yang and Shami, 2020). The hyperparameters can be determined, and the final set of variables for the regression model can be defined using the leave-one-out cross-validation method (Kudo et al., 2023). After fitting the regression model, it is crucial to analyze the errors to prevent underfitting or overfitting (Ramasubramanian and Singh, 2019). The evolution of errors in each iteration of the regression can be assessed to verify the absence of overfitting and gauge the overall quality of the final model.
@@ 2.1.1 Forest inventory of permanent plots
AGB was estimated for each tropical tree species using eq.˜1 from Calvo-Rodriguez et al. (2021):
@@ 2.5 Regression models
The modeling workflow (fig.˜6) encompassed the following:
@@ 3.1 Exploratory Analysis, Variable Selection, and Predictive Significance
The research used 2021 forest inventory data on tree height and diameter at breast height (DBH) to model the above-ground biomass ($AGB$) using equation (1). A standard Discrete Approach (DA) was applied: the total $AGB$ at each plot was calculated by summing the biomass of all individual trees, and the mean tree-level $AGB$ was then derived by dividing the total plot-level $AGB$ by the number of trees. fig.˜8 shows the values obtained from the inventory of DBH, height, and $AGB$ for the plots in this study.
@@ 3.2 Correlation analysis among variables
The correlation levels and significance of the values from the exploratory analysis are shown in figs.˜9, 10 and 11. These correlograms display the accepted forest metrics for $ALS_{D}$, $ULS_{D}$, and $SLS_{FW}$. Significant values are indicated by white “X” marks, and the correlations among the accepted variables are shown in the diagrams.
@@
fig.˜12(a) illustrates the variable importance for $AGB_{t}$ using the $ALS_{D}$ dataset. After applying a threshold of 0.5, five variables were retained: Elev.minimum, Elev.L3, Elev.MAD.mode, Elev.mode, and Elev.MAD.median, while one variable fell below this threshold and was removed. These five predictors demonstrated the strongest influence on the $AGB_{t}$ regression. For $AGB_{m}$, the relative importance of the variables is shown in fig.˜12(b). Three variables, Elev.mode, Elev.MAD.mode, and Elev.L3 surpassed the 0.5 threshold and were chosen for the $AGB_{m}$ model.
@@
fig.˜12(c) displays the variable importance for predicting $AGB_{t}$ with the $ULS_{D}$ dataset. Initially, applying the 0.5 cutoff left only one variable (Elev.skewness). To avoid underfitting, the next most influential variable (Elev.minimum) was also included, resulting in the selection of two predictors: Elev.skewness and Elev.minimum. For $AGB_{m}$ (fig.˜12(d)), three variables met the 0.5 importance threshold: Elev.maximum, Elev.minimum, and Elev.MAD.mode. Therefore, these were chosen as the final set of predictors for the $ULS_{D}$-based $AGB_{m}$ model.
@@
fig.˜12(e) shows the importance of the predictors for the simulated $SLS_{FW}$ dataset in estimating $AGB_{t}$. Strictly applying the 0.5 cutoff led to no variables meeting the criterion. Consequently, the selection rule was adjusted to include the three most influential predictors: PointDense, gLAI0t10, and lat. Finally, fig.˜12(f) depicts the importance of the variables for $AGB_{m}$ estimation using the $SLS_{FW}$ data. Here, five variables, lat, pointDense, inflGround, blairSense, and maxGround, emerged as the most significant and were thus retained for the final model.
@@ 3.4 Regression analysis
The predictive capacities of the selected models were evaluated and compared with field measurements. fig.˜13 shows that the models achieved high accuracy in predicting $AGB$.
@@ 3.4.2 Results from the SVM approach
A grid search method was employed to identify the optimal hyperparameters that minimized the estimation error. Figure 14 illustrates the error progression during this optimization phase, depicting how the adjustments to the hyperparameters impacted the model’s precision.
@@
The analysis in LABEL:tab:hyperparameters highlights the performance of SVM-based regression models across the three LiDAR platforms: ALS, ULS, and SLS. For ALS, two combinations of input variables yielded strong results. For AGBt, the best-performing variables were Elev.minimum, Elev.L3, and Elev.MAD.mode, whereas for AGBm, the most effective variables were Elev.mode and Elev.MAD.median. The estimation error (SVMr error) for AGBt was 18.0%, while for AGBm it was 13.1%, indicating that mean biomass is easier to estimate using ALS data. The optimal hyperparameters for ALS were $c=4$, $\sigma=0.25$ for AGBt, and $c=5$, $\sigma=2$ for AGBm.
@@
SVM regression was then conducted using the identified hyperparameters, and the resulting estimates were compared against the forest inventory for both AGBt and AGBm. fig.˜15 provides a synthesized view of the percentage errors, highlighting both the discrepancies and the accuracy of the model predictions.
@@ 3.5 Comparison between OLS and SVM AGB Estimates
The Mean Absolute Error (MAE) heatmap provides valuable insights into the variability of biomass estimates and their associated errors. As shown in table˜7, the differences in biomass estimates are directly correlated with the magnitude of error observed in the measurements or calculations. The upper diagonal of the table quantifies these errors, offering a clear view of how inaccuracies can propagate and lead to discrepancies in the final biomass estimates. The lower section of the same table reports the statistical significance of these variations, indicating the confidence levels associated with each comparison.
@@
As shown in Table˜8, the maximum percentage error observed in AGB modeling reached 92.7% for AGBt on plot L2.2 when utilizing ULSD data. Similarly, the most substantial modeling error for AGBm was 46.4%, recorded on the L2.2 plot using SLSFW data. In a broader context, the estimation errors for AGB were higher in L2.1 and L2.2 late-stage plots compared to L1.2, L1.1, and early and intermediate plots (E and I). The differences found in the AGB regression when using ALSD or ULSD data were relatively low, and the differences between the mean errors of ALSD and ULSD with the SLSFW data only reached 0.92% of the mean errors for AGBm and 0.68% for AGBm. The accuracy of AGB modeling using SVM exceeded 81% (ALSD: 82.0%, ULSD: 81.4%, and SLSFW: 82.9%), whereas the accuracy of OLS modeling was approximately 60%.
@@ 3.6 Results of AGB estimates for Forest Inventory, ALSD, ULSD, and SLSFW
fig.˜17 demonstrates the relationship between predicted above-ground biomass (AGB) and inventory-based AGB, expressed in Mg ha-1, across three LiDAR systems (ALSD, ULSD, and SLSFW) and different forest successional stages (Early, Intermediate, and Late). The x-axis represents the inventory-derived AGB values, considered the ground truth, whereas the y-axis displays the AGB predicted using the three LiDAR systems. A dashed diagonal line (1:1 line) is included to represent perfect agreement between the predicted and inventory-based AGB. The closer the points and trend lines are to this diagonal, the greater the prediction accuracy.